OpenStudy (jason123):

The equation of the line l is 5y-4x=10 write the standard form of the equation of the line that fits each description. A)parallel to l and passes through the pt at(-15,9). B) perpendicular to l n passes through the pt at (-15,8)

OpenStudy (radar):

Putting line 1 in slope intercept from y=(4/5)x+2 The slope is 4/5 A) A parallel line will have same slope so the equation for this line will use: 4/5=(y-9)/(x+15) cross multiplying getting 4x+60=5y-45 rearranging and combining getting: y=(4/5)x+21 y intercept form equation for the parallel line. B) Perpendicular line will have the negative reciprocal of the slope for line 1. Slope for perpendicular will be: -5/4 Equation for the perpendicular line will be: -(5/4)=(y-8)/(x+15) Cross multiplying getting: -5x-75=4y-32 4y-32=-5x-75 4y=-5x-75+32 y=-(5/4)x-43 which is the slope intercept form equation of the perpendicular line.